A solution S of sulfuric acid is prepared by dissolving 49g of H2SO4 in 500 mL of distilled water. Calculate the molar concentration of the solution


1
Expert's answer
2021-01-11T03:51:19-0500

CM=n/V

n=m/M

M(H2SO4)=98 g/mol

n(H2SO4)=49/98=0.5 mol

V=500 ml = 0.5 l

CM=0.5/0.5=1 mol/l